OpenStudy (anonymous):

The length of a rectangle is 9.7 cm more than 4 times the width. If the perimeter of the rectangle is 91.4 cm, what are its dimensions?

OpenStudy (anonymous):

perimeter = 2length + 2width = 91.4 length = (9.7 + 4w) 2(9.7+4width)+2width = 91.4 19.4 + 8width + 2width = 91.4 10 width = 72 width = 7.5 length = 9.7 + 4(7.5) length = 9.7 + 30 =39.7

OpenStudy (anonymous):

need to double check i might of made a mistake sorry